# Zooming without Zooming: Region-to-Image Distillation for Fine-Grained Multimodal Perception

### v1

@@ 1 Introduction
We achieve this by introducing a novel and easy-to-implement Region-to-Image Distillation method. Concretely, we decouple “zooming in” from inference-time (i.e., the agentic “Thinking-with-Images” process) and repurpose it as a training-time primitive. Firstly, we employ advanced models (large, cloud-hosted) to generate perception-centric questions and high-consensus answers only on micro-crops, where fine details are unambiguous and hallucinations are minimized. Then, these region-grounded data are distilled back to the full image, with explicit visual grounding via bounding-box overlays to avoid referential ambiguity. Training on such distilled supervision teaches the model (smaller, localized) to suppress global noise and pay attention to micro-level evidence directly from the full image. Therefore, the model internalizes the benefits of agentic “zooming in” into a single glance (one forward pass on the full image) at inference time without iterative tool use. This can also be summarized as an idea of “Zooming without Zooming” shown in Figure 2. Here, the first “Zooming” refers to the training-time primitive: we zoom into micro-regions to synthesize fine-grained training data and thus our model can “zoom in and out” in mind. In contrast, the second “Zooming” denotes the inference-time tool-use we seek to bypass.
@@ 3 Zooming without Zooming
Figure 3: Overview of Region-to-Image Distillation. We synthesize fine-grained VQA pairs on zoomed-in micro-crops using strong teachers with consensus filtering, then distill them to the full image via box-overlay grounding and an augmented prompt, enabling improved single-pass inference without test-time zooming.
@@ 4.2 Results
Leading Performance across Various Benchmarks. As shown in Table 2, training on our data yields consistent gains on all benchmarks, not limited to perception tasks. Notably, ZwZ-8B improves the Qwen3-VL-8B baseline from 61.52 to 68.12 average score with large gains on ZoomBench (37.87 $\rightarrow$ 58.11), and similar improvements are also observed for ZwZ-4B/7B with different scales. On the average of general perception benchmarks, ZwZ-4B and ZwZ-8B also surpass all open-source models including GLM-4.5V, Qwen3-VL-235B, and Kimi-K2.5 with much larger sizes, and remains competitive with the SOTA closed-source model Gemini-3. This clearly indicates that our data effectively distills the expertise of teacher models (GLM-4.5V and Qwen3-VL-235B) at cropped region via zooming to student models (ZwZ), and even enables the students to outperform the teachers when given the full image as global input. Beyond perception-centric benchmarks, ZwZ also reveals strong out-of-distribution performance on MMStar and BabyVision, indicating that learning to reliably retrieve micro evidence from cluttered global inputs benefits broader multimodal understanding and visual reasoning, rather than overfitting to narrow perception patterns.
@@
Outperforming Proxy Task-Based Data Synthesis Methods. Beyond comparisons with large-scale datasets, we evaluate ZwZ against methods that similarly employ synthetic data generated from proxy tasks (e.g., image jigsaw) with verifiable and reliable ground truth. These results validate that our method yield stronger, more generalizable performance, and is more practical and effective to improve small size (under 10B) models than training with proxy synthetic objectives.
@@ 6.2 Rethinking “Thinking with Images”
Information-Neutral Actions: Distill instead of Acting. In contrast, many common TwI actions concentrate on image editing (e.g., zooming (Zheng et al., 2025b), flipping (Xu et al., 2025), rotating (Zhang et al., 2025f), objection detection and segmentation (Lu et al., 2025, Zhou et al., 2025), drawing auxiliary constructs (Wei et al., 2025a, Zhao et al., 2025, Wu et al., 2025b, Wiedemer et al., 2025)) that are often predictable based on current image view and the MLLM’s strong cognition. They typically do not add new information; they mainly reduce perceptual difficulty by making existing evidence easier to access. For such cases, thinking with images at inference time can be inefficient and unnecessary, and the benefits can often be internalized into the MLLM (we assume that MLLM has enough knowledge capacity) through training with textual CoT alone, which is exactly what our method (R2I) does (Algorithm 1).
@@ 9.1 RL Training Settings
9.2 Reward Modeling in Reinforcement Learning
To provide the verifiable reward for reinforcement learning (Guo et al., 2025), we implement a multi-stage tiered reward system. This system prioritizes deterministic evaluation and falls back to heuristic or semantic methods only when necessary, ensuring a balance between accuracy and computational efficiency. The evaluation pipeline follows a three-step hierarchy:
@@ 9.2 Reward Modeling in Reinforcement Learning
Continuous Counting Reward. For tasks involving numerical counting or estimation where a binary 0/1 reward is too sparse, we employ a distance-based continuous reward for the predicted value $y_{pred}$ and the ground truth $y_{gt}$. This stage is triggered if the rule-based matching fails but both the prediction and ground truth can be parsed as numerical values. Specifically, for non-zero ground truth values, the score is defined as:
$S=\max\left(0,1-\frac{|y_{pred}-y_{gt}|}{|y_{gt}|}\right)$ (3)
In the specific case where $y_{gt}=0$, we apply a smooth exponential decay to handle the singularity:
$S=e^{-|y_{pred}-y_{gt}|}$ (4)
This continuous formulation ensures that the agent is penalized less for ”near misses,” which significantly accelerates convergence in mathematical reasoning tasks by smoothing the optimization landscape.
LLM-as-a-Judge. If neither rules nor numerical parsing can resolve the correctness, we utilize an efficient LLM (e.g., Qwen3-8B (Yang et al., 2025a)) as a judge to provide a binary 0/1 reward.
@@
Figure 8: The training reward of Qwen3-VL-8B.
9.3 Scoring System in Benchmark Evaluation
For evaluation of all the tested benchmarks, we adopt a simplified reward pipeline that mirrors the one used in reinforcement learning, but removes the continuous counting reward to keep the metric strictly verifiable and comparable across methods. In particular, we first apply rule-based matching (e.g., mathruler (Zheng, 2025)) to check exact/symbolic equivalence and assign a binary score (1 for correct, 0 otherwise). If rule-based matching fails, we fall back to an LLM-as-a-Judge using a more powerful model (e.g., Qwen3-30B (Yang et al., 2025a)) to determine correctness and output a binary 0/1 score.

### v2

@@ 1 Introduction
We achieve this by introducing a novel and easy-to-implement Region-to-Image Distillation method. Concretely, we decouple “zooming in” from inference-time (i.e., the agentic “Thinking-with-Images” process) and repurpose it as a training-time primitive. Firstly, we employ advanced teacher models (large, cloud-hosted) to generate perception-centric questions and high-consensus answers only on micro-crops, where fine details are unambiguous and hallucinations are minimized. Then, these region-grounded data are distilled back to the full image, with explicit visual grounding via bounding-box overlays to avoid referential ambiguity. Training on such distilled supervision teaches the student model (smaller, localized) to suppress global noise and pay attention to micro-level evidence directly from the full image. Therefore, the student model internalizes the benefits of agentic “zooming in” into a single glance (one forward pass on the full image) at inference time without iterative tool use. This can also be summarized as an idea of “Zooming without Zooming” shown in Figure 2. Here, the first “Zooming” refers to the training-time primitive: we zoom into micro-regions to synthesize fine-grained training data and thus our model can “zoom in and out” in mind. In contrast, the second “Zooming” denotes the inference-time tool-use we seek to bypass.
@@ 3 Zooming without Zooming
Figure 3: Overview of Region-to-Image Distillation. We synthesize fine-grained VQA pairs on zoomed-in micro-crops using strong teachers as generators, then distill them to the full image via box-overlay grounding and an augmented prompt, enabling improved single-pass inference without test-time zooming.
@@ 4.2 Results
Leading Performance across Various Benchmarks. As shown in Table 2, training on our data yields consistent gains on all benchmarks, not limited to perception tasks. Notably, ZwZ-8B improves the Qwen3-VL-8B baseline from 62.86 to 68.12 average score with large gains on ZoomBench (37.87 $\rightarrow$ 58.11), and similar improvements are also observed for ZwZ-4B/7B with different scales. On the average of general perception benchmarks, ZwZ-4B and ZwZ-8B also surpass all open-source models including GLM-4.5V, Qwen3-VL-235B, and Kimi-K2.5 with much larger sizes, and remains competitive with the SOTA closed-source model Gemini-3. This clearly indicates that our data effectively distills the expertise of teacher models (GLM-4.5V and Qwen3-VL-235B) at cropped region via zooming to student models (ZwZ), and even enables the students to outperform the teachers when given the full image as global input. Beyond perception-centric benchmarks, ZwZ also reveals strong out-of-distribution performance on MMStar and BabyVision, indicating that learning to reliably retrieve micro evidence from cluttered global inputs benefits broader multimodal understanding and visual reasoning, rather than overfitting to narrow perception patterns.
@@
Outperforming Proxy Task-Based Data Synthesis Methods. Beyond comparisons with large-scale datasets, we evaluate ZwZ against methods that similarly employ synthetic data generated from proxy tasks (e.g., image jigsaw) with verifiable and reliable ground truth. Results in Table 6 validate that our method yield stronger, more generalizable performance, and is more practical and effective to improve small size (under 10B) models than training with proxy synthetic objectives.
@@ 6.2 Rethinking “Thinking with Images”
Information-Neutral Actions: Distill instead of Acting. In contrast, many common TwI actions concentrate on image editing (e.g., zooming (Zheng et al., 2025b), flipping (Xu et al., 2025), rotating (Zhang et al., 2025f), objection detection and segmentation (Lu et al., 2025, Zhou et al., 2025), drawing auxiliary constructs (Wei et al., 2025a, Zhao et al., 2025, Wu et al., 2025b, Wiedemer et al., 2025)) that are often predictable based on current image view and the MLLM’s strong cognition. They typically do not add new information; they mainly reduce perceptual difficulty by making existing evidence easier to access. For such cases, thinking with images at inference time can be inefficient and unnecessary, and the benefits can often be internalized into the MLLM (we assume that MLLM has enough knowledge capacity) through training with textual CoT alone, which is exactly what our method (R2I) does (Algorithm 1). Besides, several prior studies (Du et al., 2025, Liu et al., 2025c) have observed that training with pure textual CoT can achieve performance comparable to agentic training when using the same training dataset (Zheng et al., 2025b, Zhang et al., 2025f). We further find that many agentic TwI models (Zheng et al., 2025b, Zhang et al., 2025f) already achieve strong performance without using any tools, which is comparable to, or even better than, using tools (Table 10). These findings also support our argument that such tool use is usually unnecessary.
Table 10: Comparison of DeepEyes and Thyme (both trained with agentic mode) under two inference modes: agentic mode (with iterative tool calls) and direct answering mode (single forward pass without any tool use). The performance gap between the two modes is surprisingly small, suggesting that the agentic overhead contributes marginally to the final accuracy on these perception benchmarks.
[…]
@@ 9.1 RL Training Settings
9.2 Reward Modeling in Reinforcement Learning and Benchmark Evaluation
To provide the verifiable reward for reinforcement learning (Guo et al., 2025) and benchmark evaluation, we implement a two-stage tiered reward system. This system prioritizes deterministic evaluation and falls back to heuristic or semantic methods only when necessary, ensuring a balance between accuracy and computational efficiency. The evaluation pipeline follows a three-step hierarchy:
@@ 9.2 Reward Modeling in Reinforcement Learning and Benchmark Evaluation
LLM-as-a-Judge. If neither rules nor numerical parsing can resolve the correctness, we utilize an efficient LLM (e.g., Qwen3-8B (Yang et al., 2025a) for RL rewarding and Qwen3-30B (Yang et al., 2025a) for benchmark evaluation) as a judge to provide a binary 0/1 reward.